# Can SSD-Mamba2 Unlock Reinforcement Learning for End-to-End Motion Control?

### v1

@@ IV-B Baseline Methods
  • Transformer Proprio-Vision. A Transformer-based model that applies self-attention over visual tokens together with a proprioceptive token, representing attention-driven cross-modal fusion [7].
@@
All baselines are trained with the same PPO algorithm, domain randomization strategy, and curriculum schedule, and are evaluated under identical experimental conditions. Where applicable, the same proprioceptive and depth encoders, along with the same token width, are reused to minimize confounding effects from differences in representation size.
IV-C Implementation Details
@@
As illustrated in Fig. 1, the network consists of lightweight encoders, an SSD-Mamba2 fusion backbone, and compact projection heads for policy and value estimation. The detailed architectural settings are summarized in Table I.
TABLE I: Network architecture settings.
@@
All policies are trained with PPO using on-policy rollouts of length $T$, minibatch updates over multiple epochs, advantage normalization, and gradient clipping. Terrain and appearance randomization are applied at episode resets, and an obstacle-density curriculum gradually increases task difficulty. The overall training procedure is outlined in Algorithm 1. Shared PPO hyperparameters are listed in Table II.
TABLE II: PPO hyperparameters.
Hyperparameter Value
Episode horizon (steps) 999
Samples per iteration 16,384
Minibatch size 1,024
Optimization epochs per update 3
Discount factor $\gamma$ 0.99
GAE parameter $\lambda$ 0.95
PPO clip parameter $\epsilon$ 0.2
Entropy coefficient 0.005
Policy learning rate $1\times 10^{-4}$
Value learning rate $1\times 10^{-4}$
Optimizer Adam
Activation function ReLU
Domain randomization.
To improve robustness and transferability, physics parameters are randomized at the beginning of each episode [44]. Unless stated otherwise, the sampled values remain fixed throughout the episode. Table III lists the randomized ranges.
TABLE III: Domain randomization ranges.
Parameter Range Units
$K_{P}$ [40, 90][40,\ 90] -
$K_{D}$ [0.4, 0.8][0.4,\ 0.8] -
Link inertia $0.5,\ 1.5]\times$ default -
Lateral friction [0.5, 1.25][0.5,\ 1.25] N·s/m
Body mass $0.8,\ 1.2]\times$ default kg
Motor friction [0.0, 0.05][0.0,\ 0.05] N·m·s/rad
Motor strength $0.8,\ 1.2]\times$ default N·m
Sensor latency [0, 0.04][0,\ 0.04] s
Visual perturbations.
Obstacle-density curriculum.
In obstacle-rich environments, training begins with sparse obstacle placement and linearly increases density to the target distribution as learning progresses [45]. This schedule is applied identically to all compared methods.
@@
IV-D Evaluation Metrics
@@
IV-E Experiment Results
IV-E1 Performance vs. Baselines
We benchmark SSD-Mamba2 against four baselines in the Thin Obstacle & Goals environment. Figure 4 shows training curves (mean $\pm$ one standard deviation across $10$ seeds), and Table IV reports final metrics.
@@ IV-E Experiment Results
TABLE IV: Performance comparison in the Thin Obstacle & Goals environment. Mean $\pm$ std over $10$ runs.
@@
From Fig. 4, SSD-Mamba2 shows a steeper early slope, higher asymptotic return, and tighter confidence band than all baselines. Numerically (Table IV), SSD-Mamba2 improves return over Transformer Proprio-Vision from $354.4$ to $537.7$ (+51.7%), reduces collisions from $202.5$ to $193.7$ (–4.3%), and increases distance from $7.55$ m to $10.50$ m (+39.1%). These results indicate both higher peak performance and more stable optimization.
IV-E2 Ablation: Effect of Multi-Modality
@@
These results confirm that multi-modality is essential for foresightful locomotion.
IV-E3 Ablation: Effect of the Fusion Backbone
@@
IV-E4 Generalization to Unseen Conditions
@@
TABLE V: Zero-shot performance on Rugged Terrain with Thin Obstacles & Goals. Mean $\pm$ std over 10 runs.
@@
TABLE VI: Zero-shot performance on Sphere Obstacles & Goals. Mean $\pm$ std over 10 runs.
@@
Mamba Vision-Only 82.74 $\pm$ 45.02 3.87 $\pm$ 0.84
@@
Rugged terrain (Table V): On uneven surfaces, proprioception alone is insufficient, yielding only $70.2$ return and $3.44$ m distance. Adding vision improves foresight: Transformer Proprio-Vision reaches $432.3$ return and $9.22$ m distance, demonstrating the benefit of multi-modal fusion. SSD-Mamba2 further improves return to $601.6$ (+39.2%) and distance to $11.31$ m (+22.7%) over Transformer Proprio-Vision, while keeping collisions competitive (638.9 vs. 560.5). The slightly higher collision rate (+14.0%) suggests that our policy explores more aggressively to achieve longer progress on rugged ground. Overall, SSD-Mamba2 attains the best performance in return and distance, underscoring the resilience of selective state-space fusion.
Sphere obstacles (Table VI): With spherical hazards, single-modality agents again fail to progress meaningfully, moving less than $6$ m on average. Transformer Proprio-Vision improves substantially ($467.6$ return, $10.26$ m distance, $163.2$ collisions). SSD-Mamba2 surpasses it with $608.5$ return (+30.1%), $12.24$ m distance (+19.3%), and slightly fewer collisions (158.8, –2.7%). This highlights that our backbone not only achieves higher asymptotic performance but also generalizes better to novel obstacle geometries.
@@
V Conclusion
@@
This work introduced SSD-Mamba2, a vision-driven cross-modal reinforcement learning framework for motion control. By encoding proprioceptive states with an MLP, patchifying depth images with a lightweight CNN, and fusing them through stacked SSD-Mamba2 layers, the method leverages state-space duality to achieve efficient long-horizon modeling with near-linear complexity. Coupled with PPO and a compact reward design under domain randomization and curriculum scheduling, the framework enables stable and scalable end-to-end training.
@@
Extensive experiments demonstrated that SSD-Mamba2 consistently outperforms proprioception-only, vision-only, and Transformer-based baselines in terms of return, safety (fewer collisions and falls), and distance traveled. Learning-curve and stability analyses further confirmed its superior convergence speed, sample efficiency, and robustness. Ablation studies highlighted the critical role of cross-modal fusion and validated the effectiveness of SSD-Mamba2 as a fusion backbone for safety-critical locomotion.
@@
While this study focused on simulated environments, future work will extend to hardware deployment, evaluating sim-to-real transfer, latency, and safety in real-world quadrupedal platforms. Beyond locomotion, the proposed approach may also benefit broader motion-control domains such as mobile navigation and human–robot interaction, where efficient fusion of multimodal signals is crucial.

### v2

@@ 4.2 Baseline Methods
  • Transformer Proprio-Vision. A Transformer-based model that applies self-attention over visual tokens together with a proprioceptive token, representing attention-driven cross-modal fusion [yang2021learning].
@@
To ensure a fair and diagnostic comparison, all baselines are trained with the same PPO algorithm, domain randomization strategy, and curriculum schedule, and are evaluated under identical experimental conditions. Wherever applicable, models share the same vision encoder, proprioceptive encoder, policy head, observation preprocessing pipeline, and training hyperparameters. This controlled setup isolates the effect of multi-modal input and the fusion and temporal modeling backbone, ensuring that observed performance differences can be attributed specifically to these factors rather than to differences in training or implementation.
4.3 Implementation Details
@@
As illustrated in Fig. 1, the network consists of lightweight encoders, an SSD-Mamba2 fusion backbone, and compact projection heads for policy and value estimation. The detailed architectural settings are summarized in Table 1.
Table 1: Network architecture settings.
@@
All policies are trained with PPO using on-policy rollouts of length $T$, minibatch updates over multiple epochs, advantage normalization, and gradient clipping. Terrain and appearance randomization are applied at episode resets, and an obstacle-density curriculum gradually increases task difficulty. The overall training procedure is outlined in Algorithm 1. Shared PPO hyperparameters are listed in Table 2.
@@
Table 2: PPO hyperparameters.
Hyperparameter Value
Episode horizon (steps) 999
Samples per iteration 16,384
Minibatch size 1,024
Optimization epochs per update 3
Discount factor $\gamma$ 0.99
GAE parameter $\lambda$ 0.95
PPO clip parameter $\epsilon$ 0.2
Entropy coefficient 0.005
Policy learning rate $1\times 10^{-4}$
Value learning rate $1\times 10^{-4}$
Optimizer Adam
Activation function ReLU
Domain randomization.
To improve robustness and transferability, physics parameters are randomized at the beginning of each episode [mehta2020active]. Unless stated otherwise, the sampled values remain fixed throughout the episode. Table 3 lists the randomized ranges.
Table 3: Domain randomization ranges.
Parameter Range Units
$K_{P}$ [40, 90][40,\ 90] -
$K_{D}$ [0.4, 0.8][0.4,\ 0.8] -
Link inertia $0.5,\ 1.5]\times$ default -
Lateral friction [0.5, 1.25][0.5,\ 1.25] N·s/m
Body mass $0.8,\ 1.2]\times$ default kg
Motor friction [0.0, 0.05][0.0,\ 0.05] N·m·s/rad
Motor strength $0.8,\ 1.2]\times$ default N·m
Sensor latency [0, 0.04][0,\ 0.04] s
Visual perturbations.
Obstacle-density curriculum.
In obstacle-rich environments, training begins with sparse obstacle placement and linearly increases density to the target distribution as learning progresses [wang2021survey]. This schedule is applied identically to all compared methods.
Parameter selection.
Parameter selection in this work follows standard practice in reinforcement learning for motion control and is kept fixed across all experiments and baselines. PPO hyperparameters and network configurations adopt commonly used settings and are not tuned per environment or per method. Task-specific parameters, including domain randomization ranges, visual perturbation strength, and obstacle-density curriculum schedules, are selected using simple and interpretable heuristics rather than exhaustive tuning. Domain randomization ranges are defined as relative perturbations around nominal physical parameters to reflect realistic uncertainty. The obstacle-density curriculum follows a monotonic schedule from sparse to dense configurations to gradually increase task difficulty. Visual perturbations are chosen to simulate typical depth-sensor artifacts without overwhelming the underlying observation structure. Overall, this design prioritizes robustness and fair comparison, ensuring that performance gains arise from the proposed framework rather than environment-specific parameter tuning.
4.4 Evaluation Metrics
@@
5 Experiment Results
5.1 Performance vs. Baselines
We benchmark SSD-Mamba2 against four baselines in the Thin Obstacle & Goals environment. Figure 4 shows training curves (mean $\pm$ one standard deviation across $10$ seeds), and Table 8 reports final metrics.
@@ 5 Experiment Results
Table 4: Performance comparison in the Thin Obstacle & Goals environment. Mean $\pm$ std over $10$ runs.
@@
From Fig. 4, SSD-Mamba2 shows a steeper early slope, higher asymptotic return, and tighter confidence band than all baselines. Numerically (Table 8), SSD-Mamba2 improves return over Transformer Proprio-Vision from $354.4$ to $537.7$ (+51.7%), reduces collisions from $202.5$ to $193.7$ (–4.3%), and increases distance from $7.55$ m to $10.50$ m (+39.1%). These results indicate both higher peak performance and more stable optimization.
5.2 Model Complexity and Resource Usage
Table 5 summarizes the model complexity and resource usage of different policy backbones, providing a quantitative view of their compute–memory trade-offs.
The Proprio-Only baseline is the lightest configuration, with 220,684220{,}684 parameters, $0.18$ M FLOPs, and minimal memory usage (14.16 MB allocated), reflecting the absence of visual processing. Introducing vision dominates the arithmetic cost: both Transformer Vision-Only and Mamba2 Vision-Only incur similar FLOPs (10.60 M vs. 10.57 M). For multi-modal fusion, SSD-Mamba2 achieves a favorable balance between model capacity and computational cost. Compared with Transformer Proprio-Vision, it reduces the parameter count from 386,476386{,}476 to 342,940342{,}940 while maintaining essentially identical FLOPs (11.01 M vs. 11.05 M). This indicates that the performance improvements reported earlier are not obtained by increasing arithmetic complexity or model size, but instead stem from more effective sequence fusion enabled by the state-space formulation.
In terms of memory usage, SSD-Mamba2 allocates and reserves more GPU memory (411.39 MB / 444.00 MB) than attention-based fusion. This overhead is primarily associated with recurrent state buffers and block-wise scan execution rather than parameter storage. Importantly, this design avoids the quadratic memory growth inherent to attention mechanisms and supports sequential, streaming-friendly computation with near-linear scaling in sequence length.
Table 5: Model complexity and resource usage
Method Total Params Model Size FLOPs Memory Allocated Memory Reserved
(MB) (M) (MB) (MB)
Proprio-Only 220,684 0.84 0.18 14.16 28.00
Transformer Vision-Only 266,092 1.02 10.60 139.57 150.00
Mamba2 Vision-Only 222,556 0.85 10.57 406.36 426.00
Transformer Proprio-Vision 386,476 1.47 11.05 167.56 188.00
SSD-Mamba2 (Ours) 342,940 1.31 11.01 411.39 444.00
5.3 Effect of Multi-Modality
@@
5.4 Effect of the Fusion Backbone
@@
5.5 Mechanism Visualization: Selectivity and Gating
To gain insight into how SSD-Mamba2 processes multimodal inputs, we visualize the internal selectivity and gating patterns for two representative layers.
Figure 5: Selectivity heatmaps of SSD-Mamba2 (Layer 1 and Layer 2). The x-axis denotes input features and the y-axis denotes selectivity channels. Brighter cells indicate stronger retention, while darker cells indicate suppression.
Selectivity. Figure 5 shows that both layers exhibit structured, non-uniform selectivity rather than uniform weighting across tokens. This indicates that the model actively learns which input channels to retain and which to suppress. Specifically, proprioceptive channels that encode stability cues consistently receive stronger weights, while subsets of depth tokens corresponding to obstacles or terrain irregularities are selectively retained. This aligns with our quantitative results: SSD-Mamba2 achieves substantially fewer collisions and longer travel distances compared with baselines, suggesting that selective filtering of proprioceptive and hazard-related cues contributes directly to safer and more foresightful locomotion.
Figure 6: Gating heatmaps of SSD-Mamba2 (Layer 1 and Layer 2). The x-axis denotes input features and the y-axis denotes gating channels. Brighter cells indicate stronger pass-through to the output, while darker cells indicate suppression.
Gating. The gating mechanism, shown in Fig. 6, regulates which processed features are propagated to the next layer after state-space scanning. We observe stronger gating for features associated with obstacle proximity and base stabilization, while redundant or noisy depth tokens are consistently suppressed. This selective pass-through acts as an information bottleneck that emphasizes task-relevant features and discards irrelevant signals. Such behavior explains two key empirical findings: (i) the faster convergence of SSD-Mamba2 in Fig. 4, since the model reduces noise early during training; and (ii) the higher asymptotic return, as policies are guided by features most critical for progress and safety.
Together, these visualizations provide a mechanistic interpretation: by selectively retaining proprioceptive stability cues, emphasizing obstacle-related depth information, and gating outputs to prioritize task-relevant features, the model achieves both efficient learning and robust control.
5.6 Generalization to Unseen Conditions
@@
Table 6: Zero-shot performance on Rugged Terrain with Thin Obstacles & Goals. Mean $\pm$ std over 10 runs.
@@
Table 7: Zero-shot performance on Sphere Obstacles & Goals. Mean $\pm$ std over 10 runs.
@@
Mamba2 Vision-Only 82.74 $\pm$ 45.02 3.87 $\pm$ 0.84
@@
Rugged terrain (Table 6): On uneven surfaces, proprioception alone is insufficient, yielding only $70.2$ return and $3.44$ m distance. Adding vision improves foresight: Transformer Proprio-Vision reaches $432.3$ return and $9.22$ m distance, demonstrating the benefit of multi-modal fusion. SSD-Mamba2 further improves return to $601.6$ (+39.2%) and distance to $11.31$ m (+22.7%) over Transformer Proprio-Vision, while keeping collisions competitive (638.9 vs. 560.5). The slightly higher collision rate (+14.0%) suggests that our policy explores more aggressively to achieve longer progress on rugged ground. Overall, SSD-Mamba2 attains the best performance in return and distance, underscoring the resilience of selective state-space fusion.
Sphere obstacles (Table 7): With spherical hazards, single-modality agents again fail to progress meaningfully, moving less than $6$ m on average. Transformer Proprio-Vision improves substantially ($467.6$ return, $10.26$ m distance, $163.2$ collisions). SSD-Mamba2 surpasses it with $608.5$ return (+30.1%), $12.24$ m distance (+19.3%), and slightly fewer collisions (158.8, –2.7%). This highlights that our backbone not only achieves higher asymptotic performance but also generalizes better to novel obstacle geometries.
@@
5.7 Impact of Input Frame Length
To examine whether longer instantaneous visual context provides additional benefits, we perform an ablation study by varying the number of stacked depth frames from 4 to 8 and 16 while keeping all other training settings unchanged. This isolates the effect of input frame length from the model’s internal temporal aggregation mechanism.
The results are summarized in Table 8. In the Thin Obstacle & Goals environment, stacking 4 frames yields the best overall performance, achieving the highest return (537.67), the fewest collisions (193.70), and the longest travel distance (10.50 m). Increasing the input length to 8 or 16 frames consistently degrades performance across all metrics, indicating that excessive visual history introduces noise rather than useful foresight in this relatively structured setting.
Table 8: Performance comparison under different input frame lengths. Mean $\pm$ std over $10$ runs.
Input Frame Length Return Collision Times Distance Moved (m)
Thin Obstacle & Goals Environment
4 537.67 $\pm$ 307.79 193.70 $\pm$ 93.61 10.50 $\pm$ 5.36
8 492.41 $\pm$ 258.61 222.53 $\pm$ 96.65 9.77 $\pm$ 3.37
16 466.72 $\pm$ 329.21 289.33 $\pm$ 145.84 8.09 $\pm$ 4.19
Rugged Terrain with Thin Obstacles & Goals Environment
4 601.62 $\pm$ 289.57 638.93 $\pm$ 118.38 11.31 $\pm$ 4.94
8 595.79 $\pm$ 324.37 583.77 $\pm$ 79.23 11.02 $\pm$ 5.29
16 391.48 $\pm$ 237.22 658.70 $\pm$ 123.47 7.69 $\pm$ 3.79
Sphere Obstacles & Goals Environment
4 608.47 $\pm$ 344.02 158.77 $\pm$ 117.62 12.24 $\pm$ 7.04
8 605.49 $\pm$ 420.51 142.10 $\pm$ 59.34 11.62 $\pm$ 6.94
16 609.38 $\pm$ 418.25 143.87 $\pm$ 93.48 11.64 $\pm$ 7.68
A similar trend is observed in the Rugged Terrain with Thin Obstacles & Goals environment. While stacking 8 frames slightly reduces collision counts (583.77 vs. 638.93), it does so at the cost of lower return and shorter travel distance. In contrast, 4-frame input achieves the highest return (601.62) and the longest distance traveled (11.31 m), suggesting a better balance between progress and safety. Stacking 16 frames leads to a pronounced performance drop, with both return and distance substantially reduced, highlighting the difficulty of exploiting long raw input sequences in uneven terrain.
In the Sphere Obstacles & Goals environment, the optimal number of input frames varies across metrics: 16 frames marginally improve return, 8 frames minimize collisions, and 4 frames maximize travel distance. This divergence further indicates that longer input sequences do not uniformly enhance control performance and may bias the policy toward specific objectives at the expense of others.
Overall, stacking 4 frames provides a favorable trade-off between perceptual context, optimization stability, and real-time responsiveness. Combined with SSD-Mamba2’s selective state-space temporal aggregation, this compact input window enables effective long-horizon reasoning while avoiding the redundancy and instability associated with longer raw input sequences.
6 Conclusion
@@
This work introduces SSD-Mamba2, a vision-driven cross-modal reinforcement learning framework for motion control. By encoding proprioceptive states with an MLP, patchifying depth images using a lightweight CNN, and fusing these representations through stacked SSD-Mamba2 layers, the method leverages state-space duality to achieve efficient long-horizon modeling with near-linear complexity. Coupled with PPO and a compact reward design under domain randomization and curriculum scheduling, the framework enables stable and scalable end-to-end training.
@@
Extensive experiments demonstrate that SSD-Mamba2 consistently outperforms proprioception-only, vision-only, and Transformer-based baselines in terms of return, safety (fewer collisions and falls), and distance traveled. Learning curves and variance across random seeds further demonstrate superior convergence speed, sample efficiency, and robustness. Ablation studies highlight the critical role of cross-modal fusion and validate the effectiveness of SSD-Mamba2 as a fusion backbone for safety-critical locomotion.
@@
While the present study focuses on empirical robustness under high-fidelity simulation, the deployment of the learned policies on physical quadrupedal platforms remains an important direction for future work. Future efforts will extend the proposed framework to real-world hardware, with particular emphasis on sim-to-real transfer through domain randomization and sensor-noise modelling, real-time inference latency under onboard computation constraints, and systematic safety validation in field conditions. In addition, the integration of explicit safety constraints and safety-aware RL mechanisms will be investigated to further enhance policy reliability during both training and execution. Beyond locomotion, the proposed approach may also benefit broader motion-control domains such as mobile navigation and human–robot interaction, where efficient fusion of multimodal signals is crucial.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment
This research was funded by the Graduate Innovation Fund under Grant 101832020CX130.
Data availability
Data will be made available on request.